#3
what does this look like in a graph after being plotted out?
First, set Y to 0 then solve for X
2X - 7
2X = -7
X = -7/2
Now plot (-7/2, 0)
Then set X to 0 and solve for Y
Y = -2(0) - 7
Y = -7
Now plot (0,-7)
1 = -2x -7
8 = -2x
8/-2 = x
x = -4
Plot (-4, 1)
I figured out this much, but cannot figure out how to get the graph. the original problem was to graph: 7.1x - 3.9y = 1.4
Help is appreciated. This is going on for quite a while and I really want to find the way to do this.

Answer by Alan3354(30959)

what does this look like in a graph after being plotted out?
First, set Y to 0 then solve for X
2X - 7
2X = -7
X = -7/2
Now plot (-7/2, 0)
Then set X to 0 and solve for Y
Y = -2(0) - 7
Y = -7
Now plot (0,-7)
1 = -2x -7
8 = -2x
8/-2 = x
x = -4
Plot (-4, 1)
I figured out this much, but cannot figure out how to get the graph. the original problem was to graph: 7.1x - 3.9y = 1.4
Help is appreciated. This is going on for quite a while and I really want to find the way to do this.
----------------------
It's a straight line, so only 2 points are needed to graph it.
7.1x - 3.9y = 1.4
Set x = 0, solve for y, as you said, but I don't know where you get your numbers.
at x = 0: 7.1*0 - 3.9y = 1.4
y = -14/39 --> (0,-14/39)
------------
at y = 0: 7.1x = 1.4
x = 14/71 --> (14/71,0)
The 2 points are (apx) (0,-0.359) and (0.197,0)